OpenStudy (anonymous):

value of arccos (sqrt2/2)... So confused!

OpenStudy (anonymous):

you just think of arccos as " the angle which the cos is ..." so the angle which the cos is \[\frac{ \sqrt{2} }{ 2}\] is 45°

OpenStudy (anonymous):

Ah, same goes for arctg and arcsin haha

OpenStudy (anonymous):

so it would be pi/4?

OpenStudy (anonymous):

yes 45° =\[\frac{ \pi }{ 4 }\]

OpenStudy (anonymous):

thanks!